A 8.47 -m ladder with a mass of 24.0 kg lies flat on the ground. A painter grabs the top end of the ladder and pulls straight upward with a force of 231 N. At the instant the top of the ladder leaves the ground, the ladder experiences an angular acceleration of 1.68 rad/s2 about an axis passing through the bottom end of the ladder. The ladder's center of gravity lies halfway between the top and bottom ends. (a) What is the net torque acting on the ladder? (b) What is the ladder's moment of inertia?
Answer #1

A)

Weight, W = mg = 24 x 9.8 = 235.2 N

Taking moment about the bottom end,

T = - (235.2 x (8.47/2)) + (231 x 8.47) = 960.5 Nm

B)

Torque, T = I a

Moment of inertia, I = T/a

I = 960.5/1.68 = 571.73 kg m^2

Comment in case any doubt please rate my answer....